OpenStudy (anonymous):

A company uses two vans to transport workers from a free parking lot to the workplace between 7:00 and 9:00 a.m. One van has 6 more seats than the other. The smaller van makes two trips every morning while the larger one makes only one trip. The two vans can transport 57 people, maximum. How many seats does the large van have

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

2x + y = 57 2x + (x + 6) = 57 3x = 51 x = 51/3 = 17 people y = x + 6 = 17 + 6 = 23 people check: 2x + y = 57 2(17) + 23 = 34 + 23 = 57 small van:17 large van:23